OpenStudy (anonymous):

how do you find the anti derivative of : 1) x(2x+1)^2 ?? please help!

OpenStudy (anonymous):

* without multiplying it all out!

OpenStudy (anonymous):

anti derivative = integration Do you know integration?

OpenStudy (inkyvoyd):

u-sub

OpenStudy (inkyvoyd):

remember the chain rule, @Laurenn_215

OpenStudy (anonymous):

yeah i know integration eccept when theres multiplying involved in integration it gets difficult!

OpenStudy (anonymous):

yeah i know the chain rule except i cant use it for this one because of the x infront?

OpenStudy (inkyvoyd):

it says, the derivative of f(g(x))=g'(x)f'(g(x))

OpenStudy (inkyvoyd):

what if we integrated both sides?

OpenStudy (inkyvoyd):

then, f(g(x))=Integration[g'(x)f'(g(x))]

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

* without multiplying it all out! --> why do you want to make your life more difficult ??

OpenStudy (inkyvoyd):

And, that's with respect to g(x)

OpenStudy (anonymous):

because when you get questions like 45x(3x - 2)^4 you dont want to spend 7000 hours in an exam multiplying it all out haha

OpenStudy (inkyvoyd):

Now, I'm going to give you a rule, and explain it.

OpenStudy (anonymous):

okay thanks inkyvoyd :)

OpenStudy (inkyvoyd):

∫ u'f(u) dx where u is a differentiable function of x =∫ f(u) du

OpenStudy (inkyvoyd):

This equation and the general method of using it for integration is called integration by u-substitution.

OpenStudy (inkyvoyd):

Why? Because we subsitute "u" for a function g(x)

OpenStudy (inkyvoyd):

Now, let me prove the rule.

OpenStudy (inkyvoyd):

*explain it

OpenStudy (anonymous):

oh! do you mean the product rule?!

OpenStudy (anonymous):

Ok. in that case The problem you just gave know can be sorted out by taking t = 3x - 2 dt = 3 dx \[\int\limits_{}^{}\frac{45}{3} \frac{t-2}{3}*t^4dt\]

OpenStudy (inkyvoyd):

∫ u'f(u) dx du/dx=u' -> rewrite in differential form du= u' dx -> put inside the indefinite integral ∫ u'f(u) dx you'll get ∫ f(u) du

OpenStudy (anonymous):

Now you can multiply it all out

OpenStudy (inkyvoyd):

This is actually an analogy of the chain rule.

OpenStudy (anonymous):

ohh

OpenStudy (inkyvoyd):

∫ u'f(u) dx, where u'=du/dx ∫ u'f(u) dx=∫ f(u)u' dx rewrite du/dx=u' in differential form -> du=u' dx look at ∫ f(u)u' dx and du= u' dx we can put du in ∫ f(u)u' dx becomes ∫ f(u) du

OpenStudy (inkyvoyd):

Now let's apply this rule to your example.

OpenStudy (inkyvoyd):

omg. Mybad. I thoguht that was x^2

OpenStudy (inkyvoyd):

That's an x?

OpenStudy (inkyvoyd):

@Laurenn_215 , my bad. This is an integration by parts problem.

OpenStudy (anonymous):

oh hehe no problem!

OpenStudy (anonymous):

@inkvoyd,not eactly

OpenStudy (inkyvoyd):

@shivam_bhalla , I can turn it into one

OpenStudy (inkyvoyd):

@Laurenn_215 , let's play around with the product rule.

OpenStudy (inkyvoyd):

the derivative of f(x)g(x)=f'(x)g(x)+f(x)g'(x) right?

OpenStudy (inkyvoyd):

Let's integrate both sides

OpenStudy (anonymous):

okay :) yep

OpenStudy (anonymous):

take t = 2x+1 dt = 2dx dt/2 = dx \[\int\limits_{}^{} \frac{t-1}{2} * t^2 * \frac{dt}{2}\] NOw multiply it out, integrate and finally substitute t back with 2x+1

OpenStudy (anonymous):

@inkyvoyd ,got it?

OpenStudy (inkyvoyd):

f(x)g(x) dx=∫ f'(x)g(x) + ∫ f(x)g'(x) dx

OpenStudy (anonymous):

yep thats the product rule inkyvod?

OpenStudy (inkyvoyd):

@shivam_bhalla , nope, but you u-subbed, so you win. I'm going to explain integrationby parts though

OpenStudy (anonymous):

Apply the chain rule and you will get the answer. \[\int\limits_{?}^{?}f(x)g(x)=f(x)\int\limits_{?}^{?}g(x)dx-\int\limits_{?}^{?}(\int\limits_{?}^{?}g(x)dx)f \prime(x)dx\] Take f(x)=x and g(x)=\[(2x+1)^{2}\] It is certain to work.

OpenStudy (anonymous):

Looks lke you want to go the long way , so continue...

OpenStudy (inkyvoyd):

f(x)g(x) dx=∫ f'(x)g(x) dx + ∫ f(x)g'(x) dx -> let's move stuff around ∫ f(x)g'(x) dx=f(x)g(x) dx-∫ f'(x)g(x) dx

OpenStudy (inkyvoyd):

set f(x) to (2x+1)^2 and g'(x) to x

OpenStudy (inkyvoyd):

then you have your answer.

OpenStudy (inkyvoyd):

Give the medal to @shivam_bhalla , because he gave a better and faster answer.

OpenStudy (inkyvoyd):

I only showed you how I would (stupidly) do it because it's a techinque that one might be interested in.

OpenStudy (anonymous):

aw okay!

OpenStudy (anonymous):

No, give it to @inkvoyd, , you will need the concept of parts later. Learnt it now

OpenStudy (anonymous):

wow guys

OpenStudy (anonymous):

*learn

OpenStudy (inkyvoyd):

uhh... thanks...

OpenStudy (anonymous):

@inkyvoyd , you are good at explaining things. A good teacher :D

OpenStudy (inkyvoyd):

Thanks :D. Unfortunately, I can't do u-sub for life lol.

OpenStudy (inkyvoyd):

Maybe I should *do* those practice problems I have lying around the house

OpenStudy (inkyvoyd):

wait, practice. NO NEVER!

OpenStudy (inkyvoyd):

*t-sub. lol

OpenStudy (anonymous):

LOL, these substitution make your life hell lot easier , so learn it. (just and advice)

OpenStudy (anonymous):

*an

OpenStudy (inkyvoyd):

I'll try ;)

OpenStudy (kropot72):

Sorry about the delay. I was well into an answer and a kind of lock up happened. Here I go again!

OpenStudy (kropot72):

The required integral is as follows: \[x/6(2x+1)^{3}-(2x+1)^{4}/48\] I will post an explanation of how it is found. Please wait.

OpenStudy (kropot72):

The formula for the integral of a product is as follows: Integral of 1st times 2nd equals1st times integral of 2nd minus integral of (differential coefficient of 1st times integral of 2nd). I will now post the working out of this.

OpenStudy (kropot72):

\[x(2x+1)^{3}/6-\int\limits_{}^{}(2x+1)^{3}/6 \]Take x to be the 1st and take (2x + 1)^2 to be the 2nd. Then using the formula we get: \[I=x(2x+1)^{3}/6-\int\limits_{}^{}(2x+1)^{3}/6\]

OpenStudy (kropot72):

Next step is to find the integral as follows: \[\int\limits_{}^{}(2x+1)^{3}/6=(2x+1)^{4}/48\]

OpenStudy (kropot72):

So we now have: \[I=x/6(2x+1)^{3}-(2x+1)^{4}/48+C\] If you differentiate the above expression you will find the original expression that was required to be integrated.

OpenStudy (kropot72):

I hope that you can follow. Please ask if you need any clarification of a step.

OpenStudy (kropot72):

The expression for the integral can be simplified down to the following: \[x ^{4}+4x ^{3}/3+x ^{2}/2+C\]